### 22 Apr 2026 v1

# A Critical Assessment of the Brain Criticality Hypothesis

Abstract
A major unresolved question in Neuroscience is: What is the origin of the observed scale-invariant correlations in neural activity? Many researchers support the “criticality hypothesis,” which proposes that the brain operates near a critical point, optimizing various information processing functions. We argue that such a critical point may not exist. Rather, the coupling between neurons and slowly varying resources (acting as “memory”), may instead generate a robust phase of neural activity with such scale-invariant correlations. This “memory-induced” long-range order (MILRO) phase is then stable to perturbations, unlike a critical point. We suggest that this MILRO phase could provide a more natural and consistent explanation of the existing experimental data than the criticality hypothesis.
I What is the Criticality Hypothesis?
The basic mechanisms of neural communication are well-established: neurons (see glossary) communicate through electrical impulses called spikes, which consume metabolic resources like Adenosine Triphosphate (ATP). At synapses—the junctions between neurons—these electrical signals trigger the release of chemical neurotransmitters that either excite or inhibit spiking in neighboring cells. This relatively simple, local interaction can generate large-scale cascades of activity throughout the brain (see Fig. 1). Yet, despite this mechanistic understanding, a fundamental question remains: how do these microscopic (and local) interactions give rise to the complex, collective dynamics observed in neural systems? The answer to this question has profound implications for understanding how the brain processes information.
Figure 1: A sketch of the human brain. Here, we consider two fundamental dynamical variables in brain behavior in a conceptual schematic: the degree of electrical activity and the amount of available resources that enable such activity. Bright yellow spikes represent bursts of electrical activity in neurons. The background colors schematically represent the degree of available resources (e.g., ATP), with a lower value indicated by red and a higher value indicated by blue. Note that the resources are most depleted in regions where neural activity has recently peaked.
One striking example of these collective dynamics appears in neuronal avalanches, the observed bursts in electrical activity when several neurons fire in quick succession. Studies have found that the probability of an avalanche of a certain size, $P(s)$, occurring (i.e., with a certain number of regions of spiking neurons involved) decays approximately as a power law with avalanche size, $P(s)\sim s^{\alpha}$, with $\alpha$ some number (see Fig. 2). [1, 5] This means that there is no single characteristic avalanche size in the brain; while small bursts of activity occur more frequently, larger ones are still expected, given sufficient time. In other words, large bursts of activity are not rare events.
To explain these experimental data, researchers have proposed the so-called criticality hypothesis[4, 21] It asserts that, during normal functioning, the brain operates in proximity of a fine-tuned critical point between two distinct phases of activity. Here, a “phase” of activity refers to a distinct mode or type of characteristic behavior (e.g., oscillations, synchronization, etc.). Physicists call points at the interface of two or more phases critical points, like the critical point at the ferromagnetic-paramagnetic phase transition. At these points, several observables decay as power laws. Hence, this could explain the detected avalanche distributions. However, as we argue below, this hypothesis may not fully explain all currently available experimental data. This, along with several other experimental and analytical considerations, has made the criticality hypothesis a source of heated debate, with many researchers in favor of [10, 2, 3] and others opposed to it. [5, 13]
Figure 2: Avalanche size, $s$, distributions extracted from electrode arrays on cultured rat cortices. Arrays of size 15, 30, and 60 were used, and a clear power-law decay in the number of activated electrodes is observed up to the size of each array. This is consistent with scale-invariant behavior over the observed range, with a cutoff set by array size. The dotted red line indicates a power-law decay of $s^{-3/2}$. Original visualization from Beggs and Plenz [1], updated version adapted from Chialvo [10].
Importantly, the criticality hypothesis does not take into account the important fact, as we have already anticipated, that neurons require resources (like ATP) to fire. As we will explain in this opinion, these resources act on much slower time scales compared to the firing of the neurons, hence providing some sort of memory to the neurons. In turn, this memory promotes a phase (not a critical point) of power-law correlations in the neural activity, thus providing a natural explanation for the observed power-law distributions of neuronal avalanche sizes.
In the remainder of this paper, we will first address some of the arguments in favor of the criticality hypothesis. Then, we will highlight a rarely discussed physical mechanism based on the important role provided by the resources, which we claim can explain much of the existing experimental data.
II Addressing the Main Arguments in Favor of The Criticality Hypothesis
II.1 A clarification on criticality, scale invariance, and long-range order
Before anything else, we must emphasize that the term “criticality” is not always used consistently. While it does have a rigorous definition in physics (as explained above), it is sometimes conflated with the notion of scale invariance. As the name implies, a system is said to be scale-invariant if it has no characteristic length scale. In other words, zooming in or out on a scale-invariant system leaves its structure unchanged.
Additionally, and almost by definition, systems with no characteristic size can have no “cut-off,” meaning power laws in some of their statistical properties are to be expected. However, although power-law distributions are hallmarks of scale invariance, they are not sufficient to prove that a system is critical. Criticality implies scale invariance [44] but the inverse is not necessarily true. There are various other ways to generate scale invariance in the absence of criticality, as several studies have amply demonstrated. [38, 32] This point is now well-recognized by proponents and detractors of the criticality hypothesis alike, but we feel that it is still worth emphasizing.
Lastly, there is the yet-weaker notion of long-range order, or long-range correlations, which is not always related to criticality. Systems with long-range order have elements that indirectly depend on, or correlate with, other elements quite strongly over long distances, much longer than their interaction length. Physicists might describe this type of system as strongly coupled. This does not necessarily mean that individual elements explicitly depend on all other elements in the systems, but rather that their implicit dependence is significant enough that it cannot be ignored. This is the case of neural activity in the brain: individual neurons are able to indirectly correlate over large distances, facilitating nuanced, macroscopic behaviors. However, as already pointed out, these long-range correlations do not necessarily imply criticality.
II.2 Properties of power-law exponents
Now that we have clarified that power laws alone are not sufficient to imply criticality, what other evidence of criticality in neural activity is there? In fact, researchers can go beyond the mere existence of power laws and probe the mathematical relationship between different power-law exponents. In particular, one can study neuronal avalanche distributions as functions of size $s$ and duration $T$ independently. Each of these distributions will decay with its own, unique power-law exponent. Additionally, one can relate average avalanche size (denoted by $\langle s\rangle$) to avalanche duration, yielding a third exponent. These relations are given below:
$P(s)\sim s^{-\alpha_{s}},\quad\qquad P(T)\sim T^{-\alpha_{T}},\quad\qquad\langle s\rangle(T)\sim T^{\gamma}.$ (1)
Sethna and other researchers [40, 39] have shown that, as long as a system is scale-invariant, the relationship between exponents $\frac{\alpha_{T}-1}{\alpha_{s}-1}=\gamma$ will hold, irrespective of whether scale invariance is related to criticality.
However, it appears that this result has been misinterpreted. Perhaps this is because it was originally provided in the context of critical systems, or because these power-law exponents are often given the moniker of “critical” exponents. Regardless, several researchers [21, 20] have used this result as a test of criticality in neural systems, even though, as we have already emphasized, scale invariance is necessary, but not sufficient, for criticality. This point has been made more explicit by the recent work of various researchers. For instance, Destexhe and Touboul, [13] Chan et al.[9] and Sun et al. [43] all explicitly showed this relationship holds in various non-critical systems.
It should also be noted that these scaling relationships enable data to be collapsed or reshaped, which can help eliminate finite-size artifacts and further verify true scale-invariance, regardless of whether or not the underlying mechanism for these power laws is criticality. [19] Again, such measures are useful for verifying the existence of scale invariance, but more sophisticated methods are required to verify whether a system is critical.
II.3 Do we really have “critical” branching in brain dynamics?
In their seminal paper, Beggs and Plenz [1] detected a power-law exponent of 3/2-3/2 in their avalanche size distributions. They also observed that heightened electrical activity in one region would, on average, lead to activity in exactly one neighboring region. Both of these results are consistent with what is known as a critical branching process, a mathematical framework which models how information is transferred in specific types of networks. The same authors then argued that this “critical” state optimizes information transfer by simultaneously enabling effective signal propagation while preventing spiking from being so prevalent that the information of where a signal originated from is washed out.
However, there is a fundamental issue with this argument. Not all detected neuronal avalanche distributions follow a power law of 3/2-3/2. In fact, several studies, both experimental [35, 23, 47, 29] and numerical [5, 34, 14, 11, 43] have observed power-law exponents which differ from 3/2-3/2. It is then unclear how relevant the precise value of these exponents is.
II.4 Point of criticality or phase of long-range order?
Perhaps the best place to look for evidence of criticality is on either side of the alleged critical point. If it truly is a critical point, then neural activity should be characteristically different on either side of this point. In other words, a small increase or decrease in model parameters away from the critical point should yield notably different behavior.
In their 2005 paper, Haldeman and Beggs [24] demonstrated that avalanche statistics did change when their cortical samples were chemically altered to encourage excitation (more spiking) or inhibition (less spiking). As expected, this yielded avalanche size distributions that deviated somewhat from power-law decays, often called “super-critical” and “sub-critical”, respectively. This has been used as evidence that the central point is critical. [2]
However, this approach suffers from a fundamental identification problem: it is unclear exactly how far from the “critical” point these chemical alterations move the system in parameter space. Without quantifying the perturbation magnitude in terms of key model parameters, the data cannot distinguish between a critical point and an extended phase of long-range order. If an entire phase of power-law distributed avalanches exists, chemical perturbations could also show the observed sensitivity to excitation/inhibition while requiring no fine-tuning whatsoever. Such a scenario—an entire phase of long-range order, robust to small perturbations and flanked by regions with heightened and diminished neural activity—would be entirely consistent with the experimental observations, yet would fundamentally contradict the criticality hypothesis.
Figure 3: A plot of correlation length alongside a set of avalanche size distributions, as reported in Sun et al. [43], obtained from modified Wilson-Cowan equations for neural activity [45, 14] solved on a square lattice of side $L$. In this visualization, the parameter $\tau_{D}$ is varied, which represents the timescale over which the available resources decay. For a wide range of $\tau_{D}$ ($25\lesssim\tau_{D}\lesssim 82$), avalanche distributions well fitted by power laws are observed, and the correlation length approaches the bounds of the lattice, both indicating scale-invariant behavior. Outside of this range, regions of minimal ($\tau_{D}\lesssim 25$) and perpetual ($\tau_{D}\gtrsim 82$) activity are observed, and the correlation length shrinks significantly. This suggests that the long-range-ordered state, which several studies have found the brain to operate at, may have been part of a wider phase of long-range order. This is difficult to reconcile with the notion that the brain must operate at or near a single critical point of a phase transition. Here, the correlation length is defined as the average distance between two units in the same avalanche. Visualization adapted from Sun et al. [43]
Surprisingly, recent research seems to suggest that this is precisely the case. Sun et al. [43] performed numerical analyses on a widely used model of neural activity [45] while gradually varying one of the model parameters. They detected a large region characterized by long-range (and even scale-invariant) avalanche distributions. On either side of this long-range-ordered phase, “down” and “up” phases exist that are characterized by minimal and perpetual neural activity, respectively. Avalanche distributions illustrating the width of this phase of long-range order can be seen in Fig. 3.
III Memory-Induced Long-Range Order
If there is indeed an entire phase of long-range order in neural activity, what generates it? Where does it come from? Recent work has indicated that the underlying cause may be the presence of memory in the system.
Here, by “memory” we do not mean information storage in the traditional sense. Rather, we are referring to time non-locality. To have time non-locality in a system, such as a collection of neurons, means that a system’s response to some stimulus depends on its entire history rather than its instantaneous state. [26] All physical systems possess some degree of time non-locality (no system can respond to a stimulus instantaneously). However, it is especially evident in systems with two types of coupled variables (degrees of freedom). Suppose one of these types evolves faster than the other. Their coupled equations might look something like this:
$\dot{\rho}$ $=F_{\rho}(\rho,R),$ (2)
$\dot{R}$ $=\frac{1}{\tau_{D}}F_{R}(\rho,R),$ (3)
where the functions $F_{\rho}$ and $F_{R}$ govern the two variables’ dynamics, and $\tau_{D}\gg 1$ is the ratio of their characteristic timescales. We can call the faster variables ($\rho$) the “primary” variables and the slower ones ($R$) the “memory” variables. This is because, assuming these types of variables depend on one another, the slower ones will act as a source of memory for the primary ones. This is precisely because they change more slowly than the other variables, so their state at a given time provides a record of the primary variables’ past dynamics.
We now consider this framework in the context of neural activity. The primary variables are the levels of electrical activity in neuronal populations, which are the easiest observables to experimentally access. These neural activities couple to several slower variables, most notably the availability of metabolic and synaptic resources (e.g., ATP, vesicle availability, and ion-gradient support) at each neuron. Indeed, many widely used models [14] of neural activity incorporate precisely these two types of variables: activities and resources.
It is clear that the amount of resources and level of electrical activity must depend on each other; neurons cannot spike without resources, and spiking will temporarily deplete available resources. Furthermore, since spiking occurs much more rapidly than the rate of growth or decay of the resources, the resources do act as a source of memory in the system. Simply by extracting the amount of resources available to a neuron at a point in time, one can infer whether or not it has been a long time since that neuron spiked. Thus, the memory variables provide a record of the neurons’ recent dynamics.
Crucially, interaction between these two types of variables does yield long-range order in the neuronal variables, and in some cases scale invariance. Sipling et al. [42] demonstrated this numerically and provided analytical arguments showing that memory in coupled variables can give rise to long-range order in the fast variables.
A few recent studies corroborate this understanding in the context of neural activity. For instance, Morrell et al. [33] demonstrated that a slowly varying external drive can lead to power-law distributions in neuronal avalanches. This slowly-varying drive acts in much the same way as an additional hidden memory degree of freedom, and can yield scale-invariant avalanche statistics over a broad parameter region rather than only at a finely tuned point. And, of course, the parameter that Sun et al. [43] vary in their work is precisely the timescale over which the resource variables decay. Only when the resources decay slowly relative to the neural activities, and thus act effectively as memories, do they observe long-range order in the neural activities.
III.1 Scope and limitations of the memory-based explanation
A key question is whether this explanation actually holds up in the face of decades of experimental work. Indeed, the emergence of long-range order due to memory is plausibly consistent with many key signatures in the existing data. The many studies that have demonstrated power-law scaling in neuronal avalanche distributions, as well as those that satisfy relationships between power-law exponents or exhibit “critical” branching, can be explained by the presence of memory-induced long-range order. Furthermore, the existence of an entire phase in which the brain can operate effectively and robustly against perturbations seems more biologically plausible than a fine-tuned point of phase transition. In addition, the criticality hypothesis would not address the recent uncovering of an entire phase of long-range order in models of neural activity. [43]
Of course, the best way to verify this behavior would be through experimentation. Specifically, since memory-induced long-range order relies upon two types of coupled variables, it would be ideal to study the degree of resource availability in the brain during normal spiking dynamics. Recent advances in genetically encoded sensors, [30] optical imaging techniques, [28, 8] and simultaneous metabolic-activity measurements [7] now make such experiments increasingly feasible. Furthermore, new research indicates that long-range order can even be “dragged” across layers in a multi-layered media. [48] Volumetric neural imaging techniques now exist [36, 12, 22] that could probe this behavior in the neocortex, which itself is organized into distinct layers. [18, 17, 25, 31]. We hope this opinion paper will spark interest in pursuing these experiments.
IV Concluding Remarks
In his 2022 paper, [3] Beggs addresses recent skepticism of the criticality hypothesis, particularly that of Destexhe and Touboul. [13] He claims that their counterexamples, which satisfy many of the criteria of critical systems without being critical themselves, are insufficient because they do not possess three key features that must exist in experimental neural systems. His point is that, in real neural systems, any signatures of criticality must:
1. Emerge due to collective behavior,
2. Possess long-range temporal correlations, and
3. Support information processing.
Unlike Destexhe and Touboul’s counterexamples, memory-induced long-range order does satisfy each of these physically motivated requirements. First, the long-range order in these systems emerges precisely due to collective behavior in neural activity, which is induced by the presence of memory as provided by the resources. Additionally, since memory is necessarily present in this system, which inherently requires variables to correlate strongly over long durations, the existence of long-range correlations in time is also expected. This has been shown explicitly in the work of Sun et al. [43] by visualizing avalanche duration distributions, which also follow power-law decays. The first two points then follow directly from the presence of memory-induced long-range order.
The notion of “supporting information processing” is more nuanced. Several studies have argued that criticality optimizes properties like information transmission, [27, 4, 37, 41] storage, [27] computational power, [6] and more. However, these studies generally emphasize the coincidence of neuronal avalanches (and other hallmarks of long-range order) with heightened performance. Crucially, they do not require criticality, only power-law correlations. As such, these properties could well be optimized in systems with long-range order that are not necessarily critical.
Additionally, it is unclear whether the “optimal” operation of several brain processing functions is even biologically necessary. For instance, Wilting and Priesemann [46] have proposed that maximizing these processing functions simultaneously may not be required. Instead, perhaps performance must merely exceed some sufficient threshold for a given task. Beyond this point, it is feasible that the more relevant biological constraint would be minimizing energy consumption.
Lastly, recent computational work has indicated that long-range order alone may be sufficient for tasks requiring high computational performance. A new computational paradigm known as memcomputing [15, 16] has demonstrated that long-range order induced by memory, when applied to solving hard computational problems, can outperform state-of-the-art algorithms at various benchmarks, supporting the claim that memory-driven collective modes can be robust without fine tuning.
Altogether, we claim that much of the recent work intending to demonstrate that the brain operates at criticality does not always distinguish between criticality and long-range order or scale invariance. Without concrete evidence for the existence of a critical point and a better understanding of the phases directly above and below it, we instead suggest that the brain may naturally operate in an entire phase of (memory-induced) long-range order. The experimental fact that resources in the brain, which promote neural activity, are always necessarily present, naturally leads to such an explanation.
V Glossary
Neuron - A nerve cell that transmits information through electrical and chemical signals.
Spike - An electrical event in a neuron characterized by a rapid increase in voltage across a neuron’s cell membrane. These electrical impulses can be transmitted between neurons, enabling communication. This event is also known as an action potential.
Synapse - The region between two neurons where chemicals produced by spikes in electrical activity can be transferred.
Neurotransmitter - A chemical released due to a spiking event that enables communication between neurons. Neurotransmitters travel across the synapse and can either promote (excite) or hinder (inhibit) the passage of electrical impulses.
Neuronal Avalanche - A collection of electrical events in neurons that occur in rapid succession. Avalanches can involve anywhere between one and all neurons in a system.
Criticality Hypothesis - The hypothesis that the brain operates near a critical point of phase transition, and that at this point, several information processing functions operate optimally due to the collective interactions between neurons. [3]
Critical Point - An individual point separating two or more distinct phases or states. For example, in a ferromagnet, the Curie temperature marks the point where spontaneous magnetization disappears and the system becomes paramagnetic.
Phase Transition - The physical transformation of a system between distinct states with unique properties, often through the variation of a single parameter (i.e., a control parameter). For example, a ferromagnet can transition to a paramagnetic state if its temperature is increased.
Memory - A general property in physical systems in which the response of a system to some stimulus depends on the full history of its past dynamics. This is also known as time non-locality [^1] In this work, we will exclusively use the word “memory” in this context, rather than in the context of the storage or retrieval of information.
Long-Range Order - The physical property characterized by a slow decay in correlation between units of a system as distance is increased. In other words, units in systems with long-range order more significantly correlate their motion with their distant neighbors than in systems without it. Equivalently, such a system could be said to have long-range correlations.
Branching Process - A mathematical model which describes how information can be transferred in networks. These processes can be further categorized by their branching parameter, which counts how many neighboring nodes an individual node transfers, on average, information to. “Critical” branching processes have a branching parameter of one.
Correlation Length - The characteristic length scale over which fluctuations in variables are correlated. Scale-invariant systems have a correlation length that diverges as the system size is increased.
VI Acknowledgments
This work was funded by the National Science Foundation via grant No. ECCS-2229880. M.D. also acknowledges funding by the Alexander von Humboldt Stiftung through the Humboldt Research Award. We thank Valérian Demaurex for useful comments on our manuscript.
VII Author Contributions
M.D. suggested and supervised the work. C. S. wrote the original draft of the manuscript. All authors contributed to the review and editing of the final version. All authors have read and approved the final manuscript.
VIII Declaration of Interests
The authors declare no competing interests.